# mode=buildmissing

### v1

@@ 5.2 Ablation studies and analysis
Total FLOPs [G $\downarrow$] 0.21 1.05 0.80 4.10
Inference Time [ms $\downarrow$] 1.2 7.0 4.5 20.1
Parameter Efficiency. Figure 7(d) demonstrates that SAQ-Decoder for toric code maintains near-constant parameter count (1.21.9M)(~1.2-1.9M) across code distances $L_{\text{code}}=4$ to $L_{\text{code}}=10$ for both noise models, exhibiting excellent scalability. In contrast, QECCT suffers from significant parameter growth, reaching $6.71M$ parameters at $L_{\text{code}}=10$ under depolarizing noise, which is a $3.5\times$ increase over our method. This difference stems from fundamental architectural choices: while QECCT processes both qubit and syndrome information through transformer layers, our approach leverages logical class prediction to decouple syndrome processing from the full qubit space dimensionality. Although both methods employ sparse attention patterns, SAQ-Decoder’s logical class embedding strategy avoids the quadratic scaling in the combined qubit-syndrome space that affects QECCT. The parameter efficiency of SAQ-Decoder directly translates to reduced memory footprint, faster training convergence, and improved deployability on quantum computing systems, making it particularly attractive for fault-tolerant implementations.
@@
Figure 7: Rotated surface code results
@@ Appendix B Constraint-Projected Nullspace Descent (CPND)
Figure 9 presents decoder recovery operator weights as a function of physical error rate $p\in\{0.05,0.10,0.15,0.20\}$ for the toric code under independent noise ($L_{\text{code}}=4$). We compare three post-processing approaches that operate on SAQ-Decoder representations: the projection baseline ($\mathbf{e}^{\prime}=\mathbf{e}^{\text{pred}}\oplus\mathbf{B}\mathbf{y}$), CPND and OSD-0 (fossorier1995soft; roffe2020decoding). OSD-0 is a post-processing algorithm that achieves minimum weight solutions while maintaining syndrome consistency, but requires matrix inversion operations that scale cubically in complexity, compared to our method’s linear complexity. Across all error rates, the methods exhibit consistent performance ordering: OSD-0 achieves the lowest weights, CPND performs comparably to OSD-0, while the projection baseline consistently yields the highest weights. The performance gaps increase monotonically with $p$, demonstrating that CPND consistently approaches minimum-weight solutions. These results highlight that structure-aware post-processing methods (CPND and OSD-0) achieve uniformly superior weight minimization compared to naive projection approaches.
@@ Appendix C Surface codes
We evaluate our method on surface codes due to their prominence in fault-tolerant quantum computing, specifically toric codes (kitaev1997quantum) which encode $k=2$ logical qubits in $n=2L_{\text{code}}^{2}$ physical qubits, and rotated surface codes (bombin2007optimal) which encode $k=1$ logical qubit in $n=L_{\text{code}}^{2}$ physical qubits. Toric codes utilize periodic boundary conditions with qubits on lattice edges, while rotated surface codes employ a lattice geometry with qubits on vertices. The stabilizer generators are organized into two distinct groups based on lattice geometry, with different implementations for each code family. For toric codes, vertex stabilizers are constructed as products of Pauli-X operators acting on all qubits adjacent to each lattice vertex, while plaquette stabilizers consist of products of Pauli-Z operators acting on qubits surrounding each lattice face, yielding a total of $m=2L_{\text{code}}^{2}-2$ stabilizer generators ($L_{\text{code}}^{2}-1$ vertex stabilizers and $L_{\text{code}}^{2}-1$ plaquette stabilizers corresponding to the vertices and faces of the $L_{\text{code}}\times L_{\text{code}}$ toric lattice). Rotated surface codes employ a fundamentally different geometry where all stabilizer generators are placed on lattice faces rather than being split between vertices and plaquettes. These face-based stabilizers come in two alternating types: X-type stabilizers (tensor products of Pauli-X operators on qubits surrounding a face) and Z-type stabilizers (tensor products of Pauli-Z operators on qubits surrounding a face). The lattice arrangement creates a natural checkerboard pattern where X- and Z-type stabilizers alternate, ensuring that every physical qubit—positioned on a vertex of the rotated lattice—participates in exactly two X-type and two Z-type stabilizer measurements. For an $L_{\text{code}}\times L_{\text{code}}$ rotated surface code, this checkerboard arrangement yields $m=2L_{\text{code}}^{2}-1$ independent stabilizer generators, with $\frac{L_{\text{code}}^{2}-1}{2}$ generators of each type, when $L$ is always odd. Figure 10 presents both code geometries.
@@
We optimize using the Adam optimizer (kingma2014adam) with mini-batches of 128512128-512 samples over 200600200-600 epochs, processing 5,00020,0005,000-20,000 mini-batches per epoch for a total of approximately $2.56\times 10^{6}$ error samples per training run. The initial learning rate is set between $3\times 10^{-4}$ and $1\times 10^{-4}$, with cosine annealing decay to $1\times 10^{-6}$ by the end of training (loshchilov2016sgdr). Detailed experimental configurations are presented in Table 2, with all experiments conducted on a 48GB NVIDIA L40S GPU.
We initialized our development from the QECCT implementation. While longer training and alternative configurations may yield further improvements, time and computational constraints limited our exploration of the hyperparameter space. We use the toric code implementation from krastanov2017deep, while the rotated surface code is implemented from scratch. Our source code is provided in the Supplementary Materials.
@@ Appendix E Comparative Threshold Analysis
This appendix provides a detailed comparison of the SAQ-Decoder’s performance threshold under depolarizing noise against recent neural decoders and high-performance classical decoders. Our numerical comparison in Table 3, the results confirm that the SAQ-Decoder’s achieved threshold of $\mathbf{18.6\%}$ approaches the theoretical Maximum Likelihood (ML) bound of $18.9\%$ (bombin2012strong) and represents the highest value reported for a practical, scalable decoder.
@@
QECCT (choukroun2024deep) 17.8%
Astra (2024) (maan2025machine) $\sim$17.0%
SU-NetQD (2025) (zhang2025self) 16.3%
ML+UF (2022) (meinerz2022scalable) 16.2%
MWPM (Classical Baseline) (wang2009threshold) 16.0%
BP-OSD (Classical Baseline) (roffe2020decoding) $\sim$16.0%
UIUF (2024) (lin2025union) 15.6%
@@ Appendix F Detailed Computational Metrics
ind QECCT 215.33 1.69 2.51e-04
11 dep SAQ 228.54 1.68 1.73e-04
dep QECCT 431.66 3.18 8.87e-04
$L$ denotes code distance. Lower values are better for all metrics.1

### v2

@@ 5.2 Ablation studies and analysis
Total FLOPs [G $\downarrow$] 0.11 0.26 0.39 0.71
Inference Time [ms $\downarrow$] 0.06 0.34 0.45 2.3
Parameter Efficiency. Figure 7(d) demonstrates that SAQ-Decoder for toric code maintains near-constant parameter count (1.21.9M)(~1.2-1.9M) across code distances $L_{\text{code}}=4$ to $L_{\text{code}}=10$ for both noise models, exhibiting excellent scalability. In contrast, QECCT suffers from significant parameter growth, reaching $6.71M$ parameters at $L_{\text{code}}=10$ under depolarizing noise, which is a $3.5\times$ increase over our method. This difference stems from fundamental architectural choices: while QECCT processes both qubit and syndrome information through transformer layers, our approach leverages logical class prediction to decouple syndrome processing from the full qubit space dimensionality. Although both methods employ sparse attention patterns, SAQ-Decoder’s logical class embedding strategy avoids the quadratic scaling in the qubit-syndrome space that affects QECCT.
@@
Figure 7: Ablation studies results.
@@ Appendix B Constraint-Projected Nullspace Descent (CPND)
Figure 9 presents decoder recovery operator weights as a function of physical error rate $p\in\{0.05,0.10,0.15,0.20\}$ for the toric code under independent noise ($L_{\text{code}}=4$). We compare three post-processing approaches that operate on SAQ-Decoder representations: the projection baseline ($\mathbf{e}^{\prime}=\mathbf{e}^{\text{pred}}\oplus\mathbf{B}\mathbf{y}$), CPND and OSD-0 (Fossorier and Lin, 1995; Roffe et al., 2020). OSD-0 is a post-processing algorithm that achieves minimum weight solutions while maintaining syndrome consistency, but requires matrix inversion operations that scale cubically in complexity, compared to our method’s linear complexity. Across all error rates, the methods exhibit consistent performance ordering: OSD-0 achieves the lowest weights, CPND performs comparably to OSD-0, while the projection baseline consistently yields the highest weights. The performance gaps increase monotonically with $p$, demonstrating that CPND consistently approaches minimum-weight solutions. These results highlight that structure-aware post-processing methods (CPND and OSD-0) achieve uniformly superior weight minimization compared to naive projection approaches.
@@ Appendix C Surface codes
We evaluate our method on surface codes due to their prominence in fault-tolerant quantum computing, specifically toric codes (Kitaev, 1997) which encode $k=2$ logical qubits in $n=2L_{\text{code}}^{2}$ physical qubits, and rotated surface codes (Bombín and Martin-Delgado, 2007) which encode $k=1$ logical qubit in $n=L_{\text{code}}^{2}$ physical qubits. Toric codes utilize periodic boundary conditions with qubits on lattice edges, while rotated surface codes employ a lattice geometry with qubits on vertices. The stabilizer generators are organized into two distinct groups based on lattice geometry, with different implementations for each code family. For toric codes, vertex stabilizers are constructed as products of Pauli-X operators acting on all qubits adjacent to each lattice vertex, while plaquette stabilizers consist of products of Pauli-Z operators acting on qubits surrounding each lattice face, yielding a total of $m=2L_{\text{code}}^{2}-2$ stabilizer generators ($L_{\text{code}}^{2}-1$ vertex stabilizers and $L_{\text{code}}^{2}-1$ plaquette stabilizers corresponding to the vertices and faces of the $L_{\text{code}}\times L_{\text{code}}$ toric lattice). Rotated surface codes employ a fundamentally different geometry where all stabilizer generators are placed on lattice faces rather than being split between vertices and plaquettes. These face-based stabilizers come in two alternating types: X-type stabilizers (tensor products of Pauli-X operators on qubits surrounding a face) and Z-type stabilizers (tensor products of Pauli-Z operators on qubits surrounding a face). The lattice arrangement creates a natural checkerboard pattern where X- and Z-type stabilizers alternate, ensuring that every physical qubit—positioned on a vertex of the rotated lattice—participates in exactly two X-type and two Z-type stabilizer measurements. For an $L_{\text{code}}\times L_{\text{code}}$ rotated surface code, this checkerboard arrangement yields $m=2L_{\text{code}}^{2}-1$ independent stabilizer generators, with $\frac{L_{\text{code}}^{2}-1}{2}$ generators of each type, when $L$ is always odd. Figure 10 presents both code geometries.
@@
We optimize using the Adam optimizer (Kingma and Ba, 2014) with mini-batches of 128512128-512 samples over 200600200-600 epochs, processing 5,00020,0005,000-20,000 mini-batches per epoch for a total of approximately $2.56\times 10^{6}$ error samples per training run. The initial learning rate is set between $3\times 10^{-4}$ and $1\times 10^{-4}$, with cosine annealing decay to $1\times 10^{-6}$ by the end of training (Loshchilov and Hutter, 2016). Detailed experimental configurations are presented in Table 2, with all experiments conducted on a 48GB NVIDIA L40S GPU.
We initialized our development from the QECCT implementation. While longer training and alternative configurations may yield further improvements, time and computational constraints limited our exploration of the hyperparameter space. We use the toric code implementation from Krastanov and Jiang (2017), while the rotated surface code is implemented from scratch. Our source code is provided in the Supplementary Materials.
@@ Appendix E Comparative Threshold Analysis
This appendix provides a detailed comparison of the SAQ-Decoder’s performance threshold under depolarizing noise against recent neural decoders and high-performance classical decoders. Our numerical comparison in Table 3, the results confirm that the SAQ-Decoder’s achieved threshold of $\mathbf{18.6\%}$ approaches the theoretical Maximum Likelihood (ML) bound of $18.9\%$ (Bombin et al., 2012) and represents the highest value reported for a practical, scalable decoder.
@@
QECCT (Choukroun and Wolf, 2024) 17.8%
Astra (2024) (Maan and Paler, 2025) $\sim$17.0%
SU-NetQD (2025) (Zhang et al., 2025) 16.3%
ML+UF (2022) (Meinerz et al., 2022) 16.2%
MWPM (Classical Baseline) (Wang et al., 2009) 16.0%
BP-OSD (Classical Baseline) (Roffe et al., 2020) $\sim$16.0%
UIUF (2024) (Lin and Lai, 2025) 15.6%
@@ Appendix F Detailed Computational Metrics
QECCT 215.33 1.69 2.51e-04
dep SAQ 228.54 1.68 1.73e-04
QECCT 431.66 3.18 8.87e-04
$L$ denotes code distance. Lower values are better for all metrics.